Circular No. 8495

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN PYXIS G. Pojmanski, Warsaw University Astronomical Observatory, reports

the discovery, via the All Sky Automated Survey (65-mm-aperture telephoto lens), of a possible nova located at α = 9^{h}18^{m}58^{s}, δ = −29^{o}42′36′′ (equinox 2000.0). ASAS V magnitudes: Mar. 5.259 UT, [14.0; 11.191, 12.0; 14.260, 12.4; 15.00, 12.6; 16.00, 12.8. H. Yamaoka, Kyushu University, reports the independent discovery by H. Haseda, Aichi, Japan, at mag 13.3 on Mar. 16.45 (0.10-m twin patrol camera), providing also the following position end figures by K. Itagaki (Teppo-cho, Yamagata, Japan, 0.60-m reflector) from his own unfiltered CCD image taken on Mar. 16.451: 58^{s}.50, 37′′.0 (red mag 12.8). Yamaoka adds that this position is close to the northern star (blue mag ∼ 18.0) of a close double on the Digitized Sky Survey.

COMET 141P/MACHHOLZ This comet has been recovered at low altitude on CCD images obtained

by M. Jäger (Stixendorf, Austria, 0.20-m Schmidt reflector; Feb. 27.8, 28.8, Mar. 1.8, and 13.8 UT) and by K. Kadota (Ageo, Japan, 0.25-m reflector; Mar. 16.4). Total visual magnitude estimates in zodiacal light: Mar. 9.10, 12.3 (A. Hale, Cloudcroft, NM, 0.41-m refl.); 11.83, 11.7 (J. J. Gonzalez, León, Spain, 0.20-m refl.). The following orbital elements, which appear with the astrometry on MPEC 2005-F08, assume this to be component A (with nongravitational parameters A_{1} = −0.27 ± 0.03, A_{2} = +0.0100 ± 0.0002; the total magnitudes in the ephemeris below assume H_{30} = 14.0).

Epoch = 2005 Mar. 11.0 TT

T = 2005 Feb. 28.24872 TT  ω = 149.^{o}28409
e = 0.7501380  Ω = 246.16135 }2000.0
q = 0.7528462 AU  i = 12.79526

a = 3.0130473 AU n^{o} = 0.18844957 P = 5.23 years

2005 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  Mag.
Mar. 11 1^{h}08.^{m}26 + 8^{o}27.^{′}3 1.486 0.775 28.^{o}5 37.^{o}8 11.5
16 1 36.93 +10 09.6 1.472 0.799 30.7 39.5 11.9
21 2 05.73 +11 40.7 1.467 0.831 32.9 40.6 12.4
26 2 34.42 +12 59.0 1.470 0.869 35.0 41.2 13.0
31 3 02.73 +14 03.6 1.483 0.912 37.1 41.3 13.7
Apr. 5 3 30.44 +14 54.2 1.504 0.959 38.9 40.9 14.3
10 3 57.32 +15 31.0 1.533 1.010 40.5 40.2 15.1
2005 March 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green